# Scaling Robust Optimization for Multi-Agent Robotic Systems: A Distributed Perspective

### v1

@@ III-A3 Robust linear chance constraints
where $\eta:\mathbb{P}_{y\in\pazocal{N}(0,1)}(y\geq q)=p$, and $\bm{\varphi}^{i}\bm{\varphi}^{i}{}^{\mathrm{T}}={\bm{\Sigma}}_{{\bm{w}}^{i}}$.
@@ III-A4 Robust covariance constraints
The covariance of the state of an agent $i$ is given as follows (derivation can be referenced from Section II of SM)
@@
It should be observed that the covariance does not depend on the deterministic uncertainty but only on the covariance of the stochastic noise. Now, let us rewrite the constraint (22) in terms of ${\bf K}^{i}$ as follows
@@
where $\bm{\varphi}^{i}\bm{\varphi}^{i}{}^{\mathrm{T}}={\bm{\Sigma}}_{{\bm{w}}^{i}}$.
@@ III-C Distributed Algorithm
Now that we have formulated the robust multi-agent trajectory optimization problem 2, we provide a distributed architecture for solving it.
@@ III-C1 Problem Setup
In this subsection, we provide an overview of steps taken to reformulate the problem 2 into a structure that would allow us to solve it in a distributed manner. We start by considering the inter-agent constraints (46) and (48). To enforce these inter-agent constraints, each agent $i\in\pazocal{V}$ needs $\{{\bf H}_{pos}{\bm{\mu}}_{x_{k},\bar{u}}^{j},{\bm{\mu}}_{d,k}^{j}\}_{k=0}^{T}$ of each of its neighbor agents $j\in\pazocal{N}_{i}$. For convenience, let us assume $\hat{{\bm{\mu}}}_{x_{k},\bar{u}}^{i}={\bf H}_{pos}{\bm{\mu}}_{x_{k},\bar{u}}^{i}$ and define the sequences $\hat{{\bm{\mu}}}_{i}^{\bar{u}}=\begin{bmatrix}\hat{{\bm{\mu}}}_{x_{0},\bar{u}}^{i};&\hat{{\bm{\mu}}}_{x_{1},\bar{u}}^{i};&\dots;\hat{{\bm{\mu}}}_{x_{T},\bar{u}}^{i}\end{bmatrix}$, ${\bm{\mu}}_{i}^{d}=\begin{bmatrix}{\bm{\mu}}_{d,0}^{i};&{\bm{\mu}}_{d,1}^{i};&\dots&{\bm{\mu}}_{d,T}^{i}\end{bmatrix}$. We can rewrite problem 2 in a simplified way as
$\{\bar{{\bm{u}}}^{i},{\bf K}^{i}\}_{i\in\pazocal{V}}=\operatornamewithlimits{argmin}$ $\;\sum_{i\in\pazocal{V}}\hat{J}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\hat{{\bm{\mu}}}_{i}^{\bar{u}},{\bm{\mu}}^{d}_{i})$
@@
$\pazocal{H}_{i,j}(\hat{{\bm{\mu}}}_{i}^{\bar{u}},\hat{{\bm{\mu}}}_{j}^{\bar{u}})\geq 0$ (57a)
$\tilde{\pazocal{H}}_{i,j}({\bm{\mu}}^{d}_{i},{\bm{\mu}}^{d}_{j})\geq 0$ (57b)
where $\hat{J}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\hat{{\bm{\mu}}}_{i}^{\bar{u}},{\bm{\mu}}^{d}_{i})=J_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i})+\pazocal{I}_{f_{i}}+\pazocal{I}_{\bar{u}^{i}}+\pazocal{I}_{{\bf K}^{i}}$
@@
To solve the problem in a distributed manner, we first introduce the copy variables $\bar{{\bm{\mu}}}_{i}^{j,\bar{u}},\bar{{\bm{\mu}}}_{i}^{j,d}$ at each agent $i$ corresponding to the variables $\hat{{\bm{\mu}}}^{\bar{u}}_{j},\bar{{\bm{\mu}}}^{d}_{j}$ of each of its neighbor agents $j\in\pazocal{N}_{i}$. Each agent then uses these copy variables to enforce the inter-agent constraints. However, these copy variables should capture the actual value of their corresponding variables. Thus, we need to enforce a consensus between the actual values of the variables $\hat{{\bm{\mu}}}_{i}^{\bar{u}},{\bm{\mu}}_{i}^{d}$ for each agent $i$ and their corresponding copy variables $\bar{{\bm{\mu}}}_{\hat{j}}^{i,\bar{u}}$, $\bar{{\bm{\mu}}}_{\hat{j}}^{i,d}$ respectively at each agent $\hat{j}$ to which agent $i$ is a neighbor of ($j\in\pazocal{P}_{i}$). Hence, the above problem can be rewritten in terms of the copy variables as follows -
$\{\bar{{\bm{u}}}^{i},{\bf K}^{i}\}_{i\in\pazocal{V}}=\operatornamewithlimits{argmin}\;\sum_{i\in\pazocal{V}}\hat{J}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\hat{{\bm{\mu}}}_{i}^{\bar{u}},{\bm{\mu}}^{d}_{i})$
@@
$\qquad\qquad\pazocal{H}_{i,j}(\hat{{\bm{\mu}}}_{i}^{\bar{u}},\bar{{\bm{\mu}}}_{i}^{j,\bar{u}})\geq 0,\;\forall\;j\in\pazocal{N}_{i},$ (58a)
$\qquad\qquad\tilde{\pazocal{H}}_{i,j}({\bm{\mu}}_{i}^{d},\bar{{\bm{\mu}}}_{i}^{j,d})\geq 0,\;\forall\;j\in\pazocal{N}_{i},$ (58b)
$\qquad\qquad\hat{{\bm{\mu}}}_{i}^{\bar{u}}=\bar{{\bm{\mu}}}_{\hat{j}}^{i,\bar{u}},\;\forall\;\hat{j}\in\pazocal{P}_{i}$ (58c)
$\qquad\qquad{\bm{\mu}}_{i}^{d}=\bar{{\bm{\mu}}}_{\hat{j}}^{i,d},\;\forall\;\hat{j}\in\pazocal{P}_{i}$ (58d)
Next, in order to use consensus ADMM (CADMM) [9], we introduce the global variables ${\bm{\nu}}_{i}^{\bar{u}},{\bm{\nu}}_{i}^{d}$ such that the constraints (58c), (58d) can be rewritten as follows.
$\hat{{\bm{\mu}}}_{i}^{\bar{u}}=\bar{{\bm{\mu}}}_{\hat{j}}^{i,\bar{u}}$ $\iff\;\hat{{\bm{\mu}}}_{i}^{\bar{u}}={\bm{\nu}}_{i}^{\bar{u}},\;{\bm{\nu}}_{i}^{\bar{u}}=\bar{{\bm{\mu}}}_{\hat{j}}^{i,\bar{u}}$ (59)
${\bm{\mu}}_{i}^{d}=\bar{{\bm{\mu}}}_{\hat{j}}^{i,d}$ $\iff\;{\bm{\mu}}_{i}^{d}={\bm{\nu}}_{i}^{d},\;\bar{{\bm{\mu}}}_{\hat{j}}^{i,d}={\bm{\nu}}_{i}^{d}$ (60)
Then, we define the local variables of each agent $i$ as $\tilde{{\bm{\mu}}}^{\bar{u}}_{i}=\begin{bmatrix}\hat{{\bm{\mu}}}_{i}^{\bar{u}};&\{\bar{{\bm{\mu}}}_{i}^{j,\bar{u}}\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$, $\tilde{{\bm{\mu}}}^{d}_{i}=\begin{bmatrix}{\bm{\mu}}_{i}^{d};&\{\bar{{\bm{\mu}}}_{i}^{j,d}\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$. For convenience, we also define the global variable sequences $\tilde{{\bm{\nu}}}_{i}^{\bar{u}}=\begin{bmatrix}{\bm{\nu}}_{i}^{\bar{u}};&\{{\bm{\nu}}_{j}^{\bar{u}}\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$, $\tilde{{\bm{\nu}}}_{i}^{d}=\begin{bmatrix}{\bm{\nu}}_{i}^{d};&\{{\bm{\nu}}_{j}^{d}\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$. Using the above-defined variables, problem (III-C1) can be rewritten as
$\{\bar{{\bm{u}}}^{i},{\bf K}^{i}\}_{i\in\pazocal{V}}$ $=\operatornamewithlimits{argmin}\;\sum_{i\in\pazocal{V}}\hat{\pazocal{J}}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\tilde{{\bm{\mu}}}^{\bar{u}}_{i},\tilde{{\bm{\mu}}}^{d}_{i})$ (61)
$\mathrm{s.t.}\;\forall\;i\in\pazocal{V},\quad\tilde{{\bm{\mu}}}^{\bar{u}}_{i}=\tilde{{\bm{\nu}}}_{i}^{\bar{u}},\quad\tilde{{\bm{\mu}}}^{d}_{i}=\tilde{{\bm{\nu}}}_{i}^{d}$
where $\hat{\pazocal{J}}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\tilde{{\bm{\mu}}}^{\bar{u}}_{i},\tilde{{\bm{\mu}}}^{d}_{i})=\hat{J}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\hat{{\bm{\mu}}}_{i}^{\bar{u}},{\bm{\mu}}^{d}_{i})+\pazocal{I}_{\pazocal{H}_{i,j}}+\pazocal{I}_{\tilde{\pazocal{H}}_{i,j}}$, and $\pazocal{I}_{\pazocal{H}_{i,j}}$, $\pazocal{I}_{\tilde{\pazocal{H}}_{i,j}}$ are indicator functions for the constraints (58a), (58b) respectively.
@@ III-C2 Distributed Approach
Now that we have reformulated the problem to the form of (61), we can use CADMM to solve it in a distributed manner. The variables $\{\tilde{{\bm{\mu}}}^{\bar{u}}_{i},\tilde{{\bm{\mu}}}^{d}_{i}\}_{i\in\pazocal{V}}$ constitute the first block, and $\{\tilde{{\bm{\nu}}}^{\bar{u}}_{i},\tilde{{\bm{\nu}}}^{d}_{i}\}_{i=1}^{N}$ constitute the second block of the CADMM. In each iteration of CADMM, the Augmented Lagragian (AL) of the problem is minimized with respect to each block in a sequential manner followed by a dual update. The AL of the problem (61) is given by
$\pazocal{L}_{\rho_{\bar{u}},\rho_{d}}=\sum_{i\in\pazocal{V}}$ $~{}\big{(}\hat{\pazocal{J}}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\tilde{{\bm{\mu}}}^{\bar{u}}_{i},\tilde{{\bm{\mu}}}^{d}_{i})$ (62)
$~{}~{}+{\bm{\lambda}}_{i}^{\bar{u}}{}^{\mathrm{T}}(\tilde{{\bm{\mu}}}^{\bar{u}}_{i}-\tilde{{\bm{\nu}}}^{\bar{u}}_{i})+{\bm{\lambda}}_{i}^{d}{}^{\mathrm{T}}(\tilde{{\bm{\mu}}}^{d}_{i}-\tilde{{\bm{\nu}}}^{d}_{i})$
$~{}~{}+\frac{\rho_{\bar{u}}}{2}||\tilde{{\bm{\mu}}}^{\bar{u}}_{i}-\tilde{{\bm{\nu}}}^{\bar{u}}_{i}||_{2}^{2}+\frac{\rho_{d}}{2}||\tilde{{\bm{\mu}}}^{d}_{i}-\tilde{{\bm{\nu}}}^{d}_{i}||_{2}^{2}\big{)}$
where ${\bm{\lambda}}_{i}^{\bar{u}}$, ${\bm{\lambda}}_{i}^{d}$ are dual variables, and $\rho_{\bar{u}}$, $\rho_{d}$ are penalty parameters.
(a) Robust case
(b) Non-robust case
Now, the update steps in an $l^{th}$ iteration of CADMM applied to the problem (61) are as follows.
@@
$\{\tilde{{\bm{\mu}}}^{\bar{u},l+1}_{i},\;\tilde{{\bm{\mu}}}^{d,l+1}_{i}\}_{i\in\pazocal{V}}$
$~{}~{}~{}=\operatornamewithlimits{argmin}\;\pazocal{L}_{\rho_{\bar{u}},\rho_{d}}(\{\tilde{{\bm{\mu}}}^{\bar{u}}_{i},\tilde{{\bm{\mu}}}^{d}_{i},\tilde{{\bm{\nu}}}^{\bar{u},l}_{i},\tilde{{\bm{\nu}}}^{d,l}_{i},{\bm{\lambda}}_{i}^{\bar{u},l},{\bm{\lambda}}_{i}^{d,l}\}_{i\in\pazocal{V}})$
Note that there is no coupling between the local variables of different agents. Thus, the above update step can be performed in a decentralized manner where each agent $i$ solves its sub-problem to update the local variables $\tilde{{\bm{\mu}}}^{\bar{u},l+1}_{i},\;\tilde{{\bm{\mu}}}^{d,l+1}_{i}$.
@@
$\{{\bm{\nu}}^{\bar{u},l+1}_{i},\;{\bm{\nu}}^{d,l+1}_{i}\}_{i\in\pazocal{V}}$
$~{}~{}~{}=\operatornamewithlimits{argmin}\;\pazocal{L}_{\rho_{\bar{u}},\rho_{d}}(\{\tilde{{\bm{\mu}}}^{\bar{u},l+1}_{i},\tilde{{\bm{\mu}}}^{d,l+1}_{i},\tilde{{\bm{\nu}}}^{\bar{u}}_{i},\tilde{{\bm{\nu}}}^{d}_{i},{\bm{\lambda}}_{i}^{\bar{u},l},{\bm{\lambda}}_{i}^{d,l}\}_{i\in\pazocal{V}})$
Let us define ${\bm{\lambda}}_{i}^{\bar{u}}$, ${\bm{\lambda}}_{i}^{d}$ as sequence of dual vectors given as ${\bm{\lambda}}_{i}^{\bar{u}}=\begin{bmatrix}\bar{{\bm{\lambda}}}_{i}^{\bar{u}};&\{\bar{{\bm{\lambda}}}_{i}^{j,\bar{u}}\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$, ${\bm{\lambda}}_{i}^{\bar{u}}=\begin{bmatrix}\bar{{\bm{\lambda}}}_{i}^{d};&\{\bar{{\bm{\lambda}}}_{i}^{j,d}\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$, such that the above update can be rewritten in a closed form for each agent $i\in\pazocal{V}$ as follows (details are provided in Section III of the SM)
${\bm{\nu}}^{\bar{u},l+1}_{i}=\frac{1}{m_{i}}\bigg{(}\frac{\bar{{\bm{\lambda}}}_{i}^{\bar{u},l}}{\rho_{\bar{u}}}+\hat{{\bm{\mu}}}_{i}^{\bar{u},l+1}+\sum_{\hat{j}\in\pazocal{P}_{i}}\big{(}\frac{\bar{{\bm{\lambda}}}_{\hat{j}}^{i,\bar{u},l}}{\rho_{\bar{u}}}+\bar{{\bm{\mu}}}_{\hat{j}}^{i,\bar{u},l+1}\big{)}\bigg{)}$ (63)
${\bm{\nu}}^{d,l+1}_{i}=\frac{1}{m_{i}}\bigg{(}\frac{\bar{{\bm{\lambda}}}_{i}^{d,l}}{\rho_{d}}+{\bm{\mu}}^{d,l+1}_{i}+\sum_{\hat{j}\in\pazocal{P}_{i}}\big{(}\frac{\bar{{\bm{\lambda}}}_{\hat{j}}^{i,d,l}}{\rho_{d}}+\bar{{\bm{\mu}}}^{i,d,l+1}_{\hat{j}}\big{)}\bigg{)}$
where $m_{i}=(n({\pazocal{P}_{i}})+1)$. After the local variables update, each agent $i$ would receive the updated variable $\bar{{\bm{\mu}}}_{\hat{j}}^{i,\bar{u},l+1}$ from all the agents $\hat{j}$ to which it is a neighbor of ($\hat{j}\in\pazocal{P}_{i}$). Using these received local variables, each agent $i$ updates the global variables ${\bm{\nu}}^{\bar{u},l+1}_{i},{\bm{\nu}}^{d,l+1}_{i}$ as per the update step (63).

### v2

@@ III-A3 Robust linear chance constraints
where $\eta:\mathbb{P}_{y\in\pazocal{N}(0,1)}(y\geq q)=p$, and $\bm{\varphi}^{i}\bm{\varphi}^{i}{}^{\mathrm{T}}={{\bm{\Sigma}}_{{\bm{w}}^{i}}}$
@@ III-A4 Robust covariance constraints
The covariance of the state of an agent $i$ is given as follows (derivation can be referenced from Appendix-B)
@@
It should be observed that the covariance does not depend on the deterministic uncertainty but only on the covariance of the stochastic noise. Now, let us rewrite the constraint (15) in terms of ${\bf K}^{i}$ as
@@
where $\bm{\varphi}^{i}\bm{\varphi}^{i}{}^{\mathrm{T}}={{\bm{\Sigma}}_{{\bm{w}}^{i}}}$
@@ III-C Distributed Algorithm
Now that we have formulated the robust multi-agent trajectory optimization problem 2, we provide a distributed architecture for solving it. The introduction of the variables $\{{\bm{\mu}}_{d,k}^{i}\}_{i\in\pazocal{V}}$ helps to characterize the state mean of the agents, which is not a point but rather a set of points. This allows us to develop a distributed framework where the agents need not share the control parameters, nor have any information of the other agents.
@@ III-C1 Problem Setup
In this subsection, we provide an overview of steps taken to reformulate the problem 2 into a structure that would allow us to solve it in a distributed manner. We start by considering the inter-agent constraints (37) and (39). To enforce these inter-agent constraints, each agent $i\in\pazocal{V}$ needs $\{{\bf H}_{pos}{\bm{\mu}}_{x_{k},\bar{u}}^{j},{\bm{\mu}}_{d,k}^{j}\}_{k=0}^{T}$ of each of its neighbor agents $j\in\pazocal{N}_{i}$. For convenience, let us assume $\hat{{\bm{\mu}}}_{x_{k},\bar{u}}^{i}={\bf H}_{pos}{\bm{\mu}}_{x_{k},\bar{u}}^{i}$ and define the sequences $\hat{{\bm{\mu}}}^{i}_{\bar{u}}=\begin{bmatrix}\hat{{\bm{\mu}}}_{x_{0},\bar{u}}^{i};&\hat{{\bm{\mu}}}_{x_{1},\bar{u}}^{i};&\dots;\hat{{\bm{\mu}}}_{x_{T},\bar{u}}^{i}\end{bmatrix}$, ${\bm{\mu}}^{i}_{d}=\begin{bmatrix}{\bm{\mu}}_{d,0}^{i};&{\bm{\mu}}_{d,1}^{i};&\dots&{\bm{\mu}}_{d,T}^{i}\end{bmatrix}$. We can rewrite problem 2 in a simplified way as
$\{\bar{{\bm{u}}}^{i},{\bf K}^{i},\hat{{\bm{\mu}}}^{i}_{\bar{u}},{\bm{\mu}}_{d}^{i}\}_{i\in\pazocal{V}}=\operatornamewithlimits{argmin}$ $\;\sum_{i\in\pazocal{V}}\hat{J}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\hat{{\bm{\mu}}}^{i}_{\bar{u}},{\bm{\mu}}_{d}^{i})$
@@
$\pazocal{H}_{i,j}(\hat{{\bm{\mu}}}^{i}_{\bar{u}},\hat{{\bm{\mu}}}^{j}_{\bar{u}})\geq 0$ (48a)
$\tilde{\pazocal{H}}_{i,j}({\bm{\mu}}_{d}^{i},{\bm{\mu}}_{d}^{j})\geq 0$ (48b)
where $\hat{J}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\hat{{\bm{\mu}}}^{i}_{\bar{u}},{\bm{\mu}}_{d}^{i})=J_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i})+\pazocal{I}_{f_{i}}+\pazocal{I}_{\bar{u}^{i}}+\pazocal{I}_{{\bf K}^{i}}$
@@
To solve the problem in a distributed manner, we first introduce the copy variables $\bar{{\bm{\mu}}}^{i}_{j,\bar{u}},\bar{{\bm{\mu}}}^{i}_{j,d}$ at each agent $i$ corresponding to the variables $\hat{{\bm{\mu}}}_{\bar{u}}^{j},\bar{{\bm{\mu}}}_{d}^{j}$ of each of its neighbor agents $j\in\pazocal{N}_{i}$. Each agent then uses these copy variables to enforce the inter-agent constraints. However, these copy variables should capture the actual value of their corresponding variables. Thus, we need to enforce a consensus between the actual values of the variables $\hat{{\bm{\mu}}}^{i}_{\bar{u}},{\bm{\mu}}^{i}_{d}$ for each agent $i$ and their corresponding copy variables $\bar{{\bm{\mu}}}^{\hat{j}}_{i,\bar{u}}$, $\bar{{\bm{\mu}}}^{\hat{j}}_{i,d}$ respectively at each agent $\hat{j}$ to which agent $i$ is a neighbor of ($\hat{j}\in\pazocal{P}_{i}$). Hence, the above problem can be rewritten in terms of the copy variables as follows -
$\{\bar{{\bm{u}}}^{i},{\bf K}^{i},\hat{{\bm{\mu}}}^{i}_{\bar{u}},{\bm{\mu}}_{d}^{i}\}_{i\in\pazocal{V}}=\operatornamewithlimits{argmin}\;\sum_{i\in\pazocal{V}}\hat{J}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\hat{{\bm{\mu}}}^{i}_{\bar{u}},{\bm{\mu}}_{d}^{i})$
@@
$\qquad\qquad\pazocal{H}_{i,j}(\hat{{\bm{\mu}}}^{i}_{\bar{u}},\bar{{\bm{\mu}}}^{i}_{j,\bar{u}})\geq 0,\;\forall\;j\in\pazocal{N}_{i},$ (49a)
$\qquad\qquad\tilde{\pazocal{H}}_{i,j}({\bm{\mu}}^{i}_{d},\bar{{\bm{\mu}}}^{i}_{j,d})\geq 0,\;\forall\;j\in\pazocal{N}_{i},$ (49b)
$\qquad\qquad\hat{{\bm{\mu}}}^{i}_{\bar{u}}=\bar{{\bm{\mu}}}^{\hat{j}}_{i,\bar{u}},\;\forall\;\hat{j}\in\pazocal{P}_{i}$ (49c)
$\qquad\qquad{\bm{\mu}}^{i}_{d}=\bar{{\bm{\mu}}}^{\hat{j}}_{i,d},\;\forall\;\hat{j}\in\pazocal{P}_{i}$ (49d)
Next, in order to use consensus ADMM (CADMM) [10], we introduce the global variables ${\bm{\nu}}^{i}_{\bar{u}},{\bm{\nu}}^{i}_{d}$ such that the constraints (49c), (49d) can be rewritten as follows.
$\hat{{\bm{\mu}}}^{i}_{\bar{u}}=\bar{{\bm{\mu}}}^{\hat{j}}_{i,\bar{u}}$ $\iff\;\hat{{\bm{\mu}}}^{i}_{\bar{u}}={\bm{\nu}}^{i}_{\bar{u}},\;{\bm{\nu}}^{i}_{\bar{u}}=\bar{{\bm{\mu}}}^{\hat{j}}_{i,\bar{u}}$ (50)
${\bm{\mu}}^{i}_{d}=\bar{{\bm{\mu}}}^{\hat{j}}_{i,d}$ $\iff\;{\bm{\mu}}^{i}_{d}={\bm{\nu}}^{i}_{d},\;\bar{{\bm{\mu}}}^{\hat{j}}_{i,d}={\bm{\nu}}^{i}_{d}$ (51)
Then, we define the local variables of each agent $i$ as $\tilde{{\bm{\mu}}}_{\bar{u}}^{i}=\begin{bmatrix}\hat{{\bm{\mu}}}^{i}_{\bar{u}};&\{\bar{{\bm{\mu}}}^{i}_{j,\bar{u}}=\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$, $\tilde{{\bm{\mu}}}_{d}^{i}=\begin{bmatrix}{\bm{\mu}}^{i}_{d};&\{\bar{{\bm{\mu}}}^{i}_{j,d}\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$. For convenience, we also define the global variable sequences $\tilde{{\bm{\nu}}}^{i}_{\bar{u}}=\begin{bmatrix}{\bm{\nu}}^{i}_{\bar{u}};&\{{\bm{\nu}}^{j}_{\bar{u}}\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$, $\tilde{{\bm{\nu}}}^{i}_{d}=\begin{bmatrix}{\bm{\nu}}^{i}_{d};&\{{\bm{\nu}}^{j}_{d}\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$. Using the above-defined variables, problem (49) can be rewritten as
$\{\bar{{\bm{u}}}^{i},{\bf K}^{i},\tilde{{\bm{\mu}}}_{\bar{u}}^{i},\tilde{{\bm{\mu}}}_{d}^{i}\}_{i\in\pazocal{V}}$ $=\operatornamewithlimits{argmin}\;\sum_{i\in\pazocal{V}}\hat{\pazocal{J}}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\tilde{{\bm{\mu}}}_{\bar{u}}^{i},\tilde{{\bm{\mu}}}_{d}^{i})$ (52)
$\mathrm{s.t.}\;\forall\;i\in\pazocal{V},\quad\tilde{{\bm{\mu}}}_{\bar{u}}^{i}=\tilde{{\bm{\nu}}}^{i}_{\bar{u}},\quad\tilde{{\bm{\mu}}}_{d}^{i}=\tilde{{\bm{\nu}}}^{i}_{d}$
where $\hat{\pazocal{J}}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\tilde{{\bm{\mu}}}_{\bar{u}}^{i},\tilde{{\bm{\mu}}}_{d}^{i})=\hat{J}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\hat{{\bm{\mu}}}^{i}_{\bar{u}},{\bm{\mu}}_{d}^{i})+\pazocal{I}_{\pazocal{H}_{i,j}}+\pazocal{I}_{\tilde{\pazocal{H}}_{i,j}}$; and $\pazocal{I}_{\pazocal{H}_{i,j}}$, $\pazocal{I}_{\tilde{\pazocal{H}}_{i,j}}$ are indicator functions for the constraints (49a), (49b) respectively.
@@ III-C2 Distributed Approach
Now that we have reformulated the problem to the form of (52), we can use CADMM to solve it in a distributed manner. The variables $\{\tilde{{\bm{\mu}}}_{\bar{u}}^{i},\tilde{{\bm{\mu}}}_{d}^{i}\}_{i\in\pazocal{V}}$ constitute the first block, and $\{\tilde{{\bm{\nu}}}_{\bar{u}}^{i},\tilde{{\bm{\nu}}}_{d}^{i}\}_{i=1}^{N}$ constitute the second block of the CADMM. In each iteration of CADMM, the Augmented Lagrangian (AL) of the problem is minimized with respect to each block in a sequential manner followed by a dual update. The AL of the problem (52) is given by
$\pazocal{L}_{\rho_{\bar{u}},\rho_{d}}=\sum_{i\in\pazocal{V}}$ $~{}\big{(}\hat{\pazocal{J}}_{i}(\bar{{\bm{u}}}^{i},{\bf K}^{i},\tilde{{\bm{\mu}}}_{\bar{u}}^{i},\tilde{{\bm{\mu}}}_{d}^{i})$ (53)
$~{}~{}+{\bm{\lambda}}^{i}_{\bar{u}}{}^{\mathrm{T}}(\tilde{{\bm{\mu}}}_{\bar{u}}^{i}-\tilde{{\bm{\nu}}}_{\bar{u}}^{i})+{\bm{\lambda}}^{i}_{d}{}^{\mathrm{T}}(\tilde{{\bm{\mu}}}_{d}^{i}-\tilde{{\bm{\nu}}}_{d}^{i})$
$~{}~{}+\frac{\rho_{\bar{u}}}{2}||\tilde{{\bm{\mu}}}_{\bar{u}}^{i}-\tilde{{\bm{\nu}}}_{\bar{u}}^{i}||_{2}^{2}+\frac{\rho_{d}}{2}||\tilde{{\bm{\mu}}}_{d}^{i}-\tilde{{\bm{\nu}}}_{d}^{i}||_{2}^{2}\big{)}$
where ${\bm{\lambda}}^{i}_{\bar{u}}$, ${\bm{\lambda}}^{i}_{d}$ are dual variables, and $\rho_{\bar{u}}$, $\rho_{d}$ are penalty parameters. Now, the update steps in an $l^{th}$ iteration of CADMM applied to the problem (52) are as follows.
@@
$\{\tilde{{\bm{\mu}}}_{\bar{u}}^{i,l+1},\;\tilde{{\bm{\mu}}}_{d}^{i,l+1}\}_{i\in\pazocal{V}}$
$~{}~{}~{}=\operatornamewithlimits{argmin}\;\pazocal{L}_{\rho_{\bar{u}},\rho_{d}}(\{\tilde{{\bm{\mu}}}_{\bar{u}}^{i},\tilde{{\bm{\mu}}}_{d}^{i},\tilde{{\bm{\nu}}}_{\bar{u}}^{i,l},\tilde{{\bm{\nu}}}_{d}^{i,l},{\bm{\lambda}}^{i,l}_{\bar{u}},{\bm{\lambda}}^{i,l}_{d}\}_{i\in\pazocal{V}})$
Note that there is no coupling between the local variables of different agents. Thus, the above update step can be performed in a decentralized manner where each agent $i$ solves its sub-problem to update the local variables $\tilde{{\bm{\mu}}}^{\bar{u},l+1}_{i},\;\tilde{{\bm{\mu}}}^{d,l+1}_{i}$ $\tilde{{\bm{\mu}}}_{\bar{u}}^{i,l+1},\;\tilde{{\bm{\mu}}}_{d}^{i,l+1}$. It is also worth noting that each agent’s local sub-problem involves constraints, such as (32) and (37), which are linearized around the previous iteration value $\tilde{{\bm{\mu}}}_{\bar{u}}^{i,l}$.
@@
$\{\tilde{{\bm{\nu}}}_{\bar{u}}^{i,l+1},\tilde{{\bm{\nu}}}_{d}^{i,l+1}\}_{i\in\pazocal{V}}$
$~{}~{}~{}=\operatornamewithlimits{argmin}\;\pazocal{L}_{\rho_{\bar{u}},\rho_{d}}(\{\tilde{{\bm{\mu}}}_{\bar{u}}^{i,l+1},\tilde{{\bm{\mu}}}_{d}^{i,l+1},\tilde{{\bm{\nu}}}_{\bar{u}}^{i},\tilde{{\bm{\nu}}}_{d}^{i},{\bm{\lambda}}^{i,l}_{\bar{u}},{\bm{\lambda}}^{i,l}_{d}\}_{i\in\pazocal{V}})$
Let us define ${\bm{\lambda}}^{i}_{\bar{u}}$, ${\bm{\lambda}}^{i}_{d}$ as sequence of dual vectors given as ${\bm{\lambda}}^{i}_{\bar{u}}=\begin{bmatrix}\bar{{\bm{\lambda}}}^{i}_{\bar{u}};&\{\bar{{\bm{\lambda}}}^{i}_{j,\bar{u}}\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$ , ${\bm{\lambda}}^{i}_{\bar{u}}=\begin{bmatrix}\bar{{\bm{\lambda}}}^{i}_{d};&\{\bar{{\bm{\lambda}}}^{i}_{j,d}\}_{j\in\pazocal{N}_{i}}\end{bmatrix}$ , such that the above update can be rewritten in a closed form for each agent $i\in\pazocal{V}$ as follows (details are provided in Appendix-I)
${\bm{\nu}}_{\bar{u}}^{i,l+1}=\frac{1}{m_{i}}\bigg{(}\frac{\bar{{\bm{\lambda}}}^{i,l}_{\bar{u}}}{\rho_{\bar{u}}}+\hat{{\bm{\mu}}}^{i,l+1}_{\bar{u}}+\sum_{\hat{j}\in\pazocal{P}_{i}}\big{(}\frac{\bar{{\bm{\lambda}}}^{\hat{j},l}_{i,\bar{u}}}{\rho_{\bar{u}}}+\bar{{\bm{\mu}}}^{\hat{j},l+1}_{i,\bar{u}}\big{)}\bigg{)}$ (54)
${\bm{\nu}}_{d}^{i,l+1}=\frac{1}{m_{i}}\bigg{(}\frac{\bar{{\bm{\lambda}}}^{i,l}_{d}}{\rho_{d}}+{\bm{\mu}}_{d}^{i,l+1}+\sum_{\hat{j}\in\pazocal{P}_{i}}\big{(}\frac{\bar{{\bm{\lambda}}}^{\hat{j},l}_{i,d}}{\rho_{d}}+\bar{{\bm{\mu}}}_{i,d}^{\hat{j},l+1}\big{)}\bigg{)}$
where $m_{i}=(n({\pazocal{P}_{i}})+1)$. After the local variables update, each agent $i$ would receive the updated variables $\bar{{\bm{\mu}}}^{\hat{j},l+1}_{i,\bar{u}}$, $\bar{{\bm{\mu}}}_{i,d}^{\hat{j},l+1}$ from all the agents $\hat{j}$ to which it is a neighbor of ($\hat{j}\in\pazocal{P}_{i}$). Using these received local variables, each agent $i$ updates the global variables ${\bm{\nu}}_{\bar{u}}^{i,l+1},{\bm{\nu}}_{d}^{i,l+1}$ as per the update step (54) .